# Numerical approximation of McKean-Vlasov SDEs via stochastic gradient descent

### v2

@@ Abstract
We propose a novel approach to numerically approximate McKean-Vlasov stochastic differential equations (MV-SDE) using stochastic gradient descent (SGD) while avoiding the use of interacting particle systems. The SGD technique is deployed to solve a Euclidean minimization problem, obtained by first representing the MV-SDE as a minimization problem over the set of continuous functions of time, and then approximating the domain with a finite-dimensional subspace. Convergence is established by proving certain intermediate stability and moment estimates of the relevant stochastic processes, including the tangent processes. Numerical experiments illustrate the competitive performance of our SGD based method compared to the IPS benchmarks. This work offers a theoretical foundation for using the SGD method in the context of numerical approximation of MV-SDEs, and provides analytical tools to study its stability and convergence.
@@ 1 Introduction
MV-SDEs play an important role in a variety of fields, including, but not limited to, kinetic theory of gases and plasmas [13, 6], chemotaxis modelling [38], population dynamics including crowd [33] and pedestrian [32] dynamics, Machine Learning (ML) [11, 12], finance and optimal control [10], quantum mechanics [23], and mean-field games [10]. Typically, the solution of an MV-SDE is approximated via an $N_{0}$-dimensional interacting particle system (IPS) and it is well-known, via the so-called “propagation of chaos” results (e.g. [39, 34]), that the empirical law of the interacting particle system converges to the law of the MV-SDE as $N_{0}\to\infty$. Numerically, one can then use time-discretization and simulation to obtain accurate approximate solutions of the MV-SDE via the IPS ([9, 2, 41]). The applicability of this approach, has been recently extended across a variety of settings under varying regularity and growth assumptions on its coefficients, see for example, [24], [37], [16],[17], [8].
@@
Even though the “IPS” approach is very powerful and can be applied to many different classes of MV-SDEs, depending on the type of interaction kernel, the computational cost of simulating an interacting particle system can be expensive. As such, there have been many other works in the literature to explore alternate techniques for numerically solving for the solution of an MV-SDE. In [40], an iterative multilevel Monte Carlo scheme of creating a particle system is combined with Picard iteration to reduce the computational cost. In [21], the technique of analytical expansions is used to derive accurate approximation of the density of the solution of an MV-SDE with elliptic constant diffusion coefficient and bounded drift. In another approach, [3], [22] use the Fokker-Planck equation representation of the density of the solution to apply numerical techniques used to solve partial differential equations (PDEs). In [1], the authors consider a class of MV-SDEs for which the solution can be readily computed using the Picard iteration. They use a Fourier transform approach to solve a fixed point equation, which also allowed them to incorporate Lévy jumps in their analysis in a relatively straightforward manner. In [28], the authors used a multilevel Picard iteration to develop numerical estimates of the solution of a class of MV-SDEs in high dimensions. They are able to prove that their numerical approach is computationally efficient and avoids the curse of dimensionality unlike the IPS approach. In [14] the authors developed a cubature method based approach to solve decoupled forward-backward MV-SDEs.
@@
Recently, there have been tremendous developments in using machine learning techniques to efficiently solve nonlinear PDEs in high dimensions owing to the seminal work [25]. Typically in this approach, an appropriate approximation of the forward process related to the nonlinear PDE and an approximation of the solution in terms of its gradient are provided as inputs to a feedforward neural network. The network is then trained to learn the mapping between the forward process and the gradient of the solution using a SGD method which minimizes a suitable loss function based on the terminal condition of the nonlinear PDE. It has been proven under different settings of nonlinear PDEs (for example, [27], [26]) that the machine-learning-based method reliant on SGD avoids the curse of dimensionality. This approach of using machine learning to estimate the solution of a nonlinear PDE has also been successfully extended to solve McKean-Vlasov stochastic control problems, see, for example, [11], [12]. Our proposed approach to solve an MV-SDE using SGD is inspired by these aforementioned works (and the references therein). We do not rely on a neural network to learn the solution but instead use SGD to learn optimal weights in a polynomial basis function representation of the unknown functions in the considered MV-SDE.
@@
Overall, the results of this paper can be viewed as a proof of concept for the application of stochastic gradient descent algorithms as a valid tool to deliver numerical approximations of MV-SDEs without the simulation of IPSs. In order to reduce the technicalities and keep the proofs as transparent as possible, the convergence analysis is performed here in the case of bounded coefficients. However, it could be extended to the case of $\alpha_{j}(t,\cdot)$, $\beta_{j}(t,\cdot)$ with sub-linear growth and $\varphi_{1,j}$ and $\varphi_{2,j}$ with polynomial growth. We discuss this issue further below. In particular, the numerical tests presented in Section 4.2 confirm that the approach is viable also when the coefficients $\varphi_{1,j}$ and $\varphi_{2,j}$ have quadratic growth. An extension of this methodology to the case of Hölder continuous drift coefficients seems feasible, up to dealing with some technical issues arising from the fact that the gradient process becomes a diffusion with singular (distributional) drift coefficients. We also expect this approach to be suitable for extensions to the jump-diffusion framework. Finally, an interesting and challenging extension is in regards the adapting this methodology to frameworks where the simulation of the associated IPS is less amenable, e.g. conditional MV-SDEs or MV-SDEs with moderate interaction among others.
@@
$L^{p}:=L^{p}([0,T],\mathbb{R}^{m})$ Space of Borel-measurable functions $f:[0,T]\to\mathbb{R}^{m}$ with
[…]
$L^{\infty}:=L^{\infty}([0,T],\mathbb{R}^{m})$ Space of Borel-measurable functions $f:[0,T]\to\mathbb{R}^{m}$ with
$\|f\|_{L^{\infty}}=\sup_{t\in[0,T]}|f(t)|<\infty,$ for $p\geq 1$
@@
$C^{n}_{b}([0,T])$ Space of $n$-times continuously differentiable functions $f:[0,T]\to\mathbb{R}^{m}$ with
bounded derivatives
@@
$L^{p}$ Space of random variables $\xi$ such that $\mathbb{E}[|\xi|^{p}]<\infty,$ for $p\geq 1$
[…]
for $p\geq 1$
$\mathcal{S}^{p}_{T}$ Space of progressively measurable processes $X$ such that
$\mathbb{E}[\sup_{t\in[0,T]}|X_{t}|^{p}]<\infty,$ for $p\geq 1$
$\mathcal{S}^{\infty}_{T}$ Space of progressively measurable processes $X$ such that
$\operatorname*{ess\,sup}_{\omega\in\Omega}\sup_{0\leq t\leq T}|X_{t}|<\infty$
@@
$\mathscr{L}$ Lifting map (19), $\mathbb{R}^{(n+1)\times K}\ni a\mapsto\mathscr{L}a:=a_{0}g_{0}+\cdots+a_{n}g_{n}\in L^{\infty}$, with
$g_{0},\cdots,g_{n}$ being linearly independent vectors of $C([0,T],\mathbb{R})$
@@ 2 McKean-Vlasov SDEs with separable coefficients
Let $T>0$ be a fixed time horizon. All the stochastic differential equations in this article are considered up to time $T$. We consider the class of $\mathbb{R}^{d}$-valued MV-SDEs in (1), where the interaction through the state is separated from the interaction through the law. Equivalently, their dynamics can be written in a more compactly as
@@
for measurable maps $\alpha:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{K\times d}$, $\beta:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{K\times d\times q}$, $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{K}$, and with $W$ being a $q$-dimensional Brownian motion (column vector).
@@
The above result follows from [10, Theorem 4.21 (p.236)]). In particular, the fact that $\bar{\gamma}$ is bounded and continuous is straightforward as $\varphi$ is bounded and continuous.
@@
Concerning the well-posedness of MV-SDEs and the moment estimates of its solutions, Lipschitz continuity of the coefficients is a rather standard assumption. The literature on MV-SDEs also contains results for less regular coefficients (see [15] and the references therein), but these typically require additional structural assumptions, e.g. ellipticity conditions for the diffusion coefficient. Also, we rely on the Lipschitz continuity of $\alpha(t,\cdot)$ and $\beta(t,\cdot)$ in the convergence analysis of the SGD algorithm (see Section 1 below). Thus, we prefer to introduce this regularity requirement already at this stage.
@@
$[0,T]\times\mathbb{R}^{d}\ni(t,x)\mapsto\big{(}\mathscr{K}^{m}u\big{)}(t,x)=\underbrace{\mathscr{K}\big{(}\cdots\mathscr{K}(\mathscr{K}u)\cdots\big{)}}_{m\text{ times}}(t,x).$ (6)
@@ 2.1 Fixed-point equation for $\bar{\gamma}$
$F:C\big{(}[0,T],\mathbb{R}^{K}\big{)}\to\mathbb{R},\qquad F(\gamma):=\|\Psi(\gamma)-\gamma\|_{L^{2}([0,T])}.$ (12)
@@ 2.2 Projection on a finite dimensional domain
Finally note that, in particular, the distance $d_{n}(\gamma)$ and the projection $p_{n}(\gamma)$ (see Remark 2.15) are the same for all these choices of basis, for they all generate the same finite dimensional space $S_{n}=\mathcal{P}_{n}(\mathcal{R}^{K})$. In this case, $p_{n}$ in (20) is called the polynomial of best approximation of $\gamma$ (see [29, Chapter 5]).
@@
Since the maps $\gamma$ are bounded by $\|\varphi\|_{\infty}$, we introduce the auxiliary map ${\bf h}:\mathbb{R}^{K}\to\mathbb{R}^{K}$ so that ${\bf h}(\mathscr{L}a)$ is also bounded. Concretely, let ${\bf h}$ be a smooth bounded function $\mathbb{R}^{K}\ni x\mapsto{\bf h}(x)=\big{(}{\bf h}_{1}(x_{1}),\cdots,{\bf h}_{K}(x_{K})\big{)}$ defined as
@@
The boundedness of $\varphi$ implies that $F^{2}({\bf h}\circ\mathscr{L}a)$ is bounded. Thus, by (25), $G(a)$ tends to +$\infty$+ ∞ as $|a|\to\infty$. On the other hand, Lemma 2.13 together with the continuity of ${\bf h}$ and $H$ imply that $G$ is continuous on $\mathbb{R}^{(n+1)K}$. Therefore, the fact that $G$ has at least a minimum $\bar{a}_{n}$ on $\mathbb{R}^{(n+1)K}$ stems from Weierstrass’ extreme value theorem.
@@
A theorem of D. Jackson (see [42, Section 5.1.5 (p.261)]) then yields
@@ 3.1 Some preliminaries
    Let now $W$, $\tilde{W}$ be two independent Brownian motions and $\xi$, $\tilde{\xi}$ be two identically distributed independent random variables, which are also independent of $(W,\tilde{W})$. Assuming weak-uniqueness for the solutions to (38) we have
@@ 3.2 Algorithm
    • $(W_{m})_{m\in\mathbb{N}}$ and $(\tilde{W}_{m})_{m\in\mathbb{N}}$ are sequences of independent $q$-dimensional standard Brownian motions adapted to (¯$(\bar{\mathcal{F}}_{m})_{m\in\mathbb{N}}$. Namely, for any $m\in\mathbb{N}$, $W_{m}=(W_{m,t})_{t\in[0,T]}$ is a $q$-dimensional standard Brownian motion such that $\mathcal{F}^{W_{m}}_{T}\subset\bar{\mathcal{F}}_{m}$, and the same holds for $\tilde{W}_{m}$;
@@
Input : Maximum number of iterations $m_{\max}$;

### v3

@@ Abstract
We propose a novel approach to numerically approximate McKean-Vlasov stochastic differential equations (MV-SDE) using stochastic gradient descent (SGD) while avoiding the use of interacting particle systems (IPS) and the associated simulation costs required to achieve the “propagation of chaos” limit. The SGD technique is deployed to solve a Euclidean minimization problem, obtained by first representing the MV-SDE as a minimization problem over the set of continuous functions of time, and then approximating the domain with a finite-dimensional subspace. Convergence is established by proving certain intermediate stability and moment estimates of the relevant stochastic processes, including the tangent processes. Numerical experiments illustrate the competitive performance of our SGD based method compared to the IPS benchmarks. This work offers a theoretical foundation for using the SGD method in the context of numerical approximation of MV-SDEs, and provides analytical tools to study its stability and convergence.
@@ 1 Introduction
MV-SDEs play an important role in a variety of fields, including, but not limited to, kinetic theory of gases and plasmas [12, 38], chemotaxis modelling [45], population dynamics including crowd [39] and pedestrian [37] dynamics, machine learning (ML) [10, 11], finance and optimal control [9], quantum mechanics [27], and mean-field games [9]. Typically, the solution of an MV-SDE is approximated via an $N_{0}$-dimensional interacting particle system (IPS) and it is well-known, via the so-called “propagation of chaos” results (e.g. [46, 40]), that the empirical law of the interacting particle system converges to the law of the MV-SDE as $N_{0}\to\infty$. Numerically, one can then use time-discretization and simulation to obtain accurate approximate solutions of the MV-SDE via the IPS ([8, 2, 48]). The applicability of this approach, has been recently extended across a variety of settings under varying regularity and growth assumptions on its coefficients, see for example, [28], [44], [17],[18], [7], [19], [20], [50], [42].
@@
Even though the “IPS” approach is very powerful and can be applied to many different classes of MV-SDEs, depending on the type of interaction kernel, the computational cost of simulating an interacting particle system can be expensive. As such, there have been many other works in the literature to explore alternate techniques for numerically solving for the solution of an MV-SDE. In [47], an iterative multilevel Monte Carlo scheme of creating a particle system is combined with Picard iteration to reduce the computational cost. In [25], the technique of analytical expansions is used to derive accurate approximation of the density of the solution of an MV-SDE with elliptic constant diffusion coefficient and bounded drift. In another approach, [3], [26], [13] use the Fokker-Planck equation representation of the density of the solution to apply numerical techniques used to solve partial differential equations (PDEs). In [1], the authors consider a class of MV-SDEs for which the solution can be readily computed using the Picard iteration. They use a Fourier transform approach to solve a fixed point equation, which also allowed them to incorporate Lévy jumps in their analysis in a relatively straightforward manner. In [32], the authors used a multilevel Picard iteration to develop numerical estimates of the solution of a class of MV-SDEs in high dimensions. They are able to prove that their numerical approach is computationally efficient and avoids the curse of dimensionality unlike the IPS approach. In a way, many of the methods mentioned are of decoupling-type: in a first step the measure component (or a functional thereof) is approximated, said approximation is plugged back into the Mckean-Vlasov equation’s dynamics yielding an SDE that is then approximated in the usual fashion—such approaches have been used for importance sampling [21] or to deploy Koopman transfer operator approaches to the non-linear semi-group operator of MV-SDEs [33]. In [15] the authors developed a cubature method based approach to solve decoupled forward-backward MV-SDEs. Noteworthy of mention is the recent work of [14] in the context of ergodic simulation of the invariant distribution of McKean-Vlasov equations—a setting we do not explore. There, one replaces the measure argument in the coefficients of the MV-SDE by the process’ occupation (empirical) measure. This transforms the MV-SDE into a self‐interacting diffusion that has full history dependence in its coefficients. The methodology remains then single‐trajectory (no interacting particles) and thus avoiding the classical particle‐system propagation‐of‐chaos costs.
@@
Recently, there have been tremendous developments in using machine learning techniques to efficiently solve nonlinear PDEs in high dimensions o wing to the seminal work [29]. Typically in this approach, an appropriate approximation of the forward process related to the nonlinear PDE and an approximation of the solution in terms of its gradient are provided as inputs to a feedforward neural network. The network is then trained to learn the mapping between the forward process and the gradient of the solution using a SGD method which minimizes a suitable loss function based on the terminal condition of the nonlinear PDE. It has been proven under different settings of nonlinear PDEs (for example, [31], [30]) that the machine-learning-based method reliant on SGD avoids the curse of dimensionality. This approach of using machine learning to estimate the solution of a nonlinear PDE has also been successfully extended to solve McKean-Vlasov stochastic control problems, see, for example, [10], [11]. Our proposed approach to solve an MV-SDE using SGD is inspired by these aforementioned works (and the references therein). We do not rely on a neural network to learn the solution but instead use SGD to learn optimal weights in a polynomial basis function representation of the unknown functions in the considered MV-SDE. The particular class of MV-SDE we consider in this work (2) allows us to develop a more targeted approach, which is naturally more efficient than a general neural network based estimation approach.
@@
Overall, the results of this paper can be viewed as a proof of concept for the application of stochastic gradient descent algorithms as a valid tool to deliver numerical approximations of MV-SDEs without the simulation of IPSs. In order to reduce the technicalities and keep the proofs as transparent as possible, the convergence analysis is performed here in the case of bounded coefficients. However, it could be extended to the case of $\alpha_{j}(t,\cdot)$, $\beta_{j}(t,\cdot)$ with sub-linear growth and $\varphi_{1,j}$ and $\varphi_{2,j}$ with polynomial growth. We discuss this issue further below (see Remark 2.5). In particular, the numerical tests presented in Section 4.2 confirm that the approach is viable also when the coefficients $\varphi_{1,j}$ and $\varphi_{2,j}$ have quadratic growth. An extension of this methodology to the case of Hölder continuous drift coefficients seems feasible, up to dealing with some technical issues arising from the fact that the gradient process becomes a diffusion with singular (distributional) drift coefficients. We also expect this approach to be suitable for extensions to the jump-diffusion framework. Finally, an interesting and challenging extension is in regards the adapting this methodology to frameworks where the simulation of the associated IPS is less amenable, e.g. conditional MV-SDEs or MV-SDEs with moderate interaction among others.
@@
$L^{p}([0,T],\mathbb{R}^{m})$ (or $L^{p}([0,T])$ in short) is the Space of Borel-measurable functions
$f:[0,T]\to\mathbb{R}^{m}$ with $\|f\|_{L^{p}([0,T])}=\left(\int_{0}^{T}|f(t)|^{p}dt\right)^{\frac{1}{p}}<\infty,$ for $p\geq 1$.
$L^{\infty}([0,T],\mathbb{R}^{m})$ (or $L^{\infty}([0,T])$ in short) is the Space of Borel-measurable functions
$f:[0,T]\to\mathbb{R}^{m}$ with $\|f\|_{L^{\infty}([0,T])}=\sup_{t\in[0,T]}|f(t)|<\infty$
@@
$C^{n}_{b}([0,T])$ Space of $n$-times continuously differentiable functions $f:[0,T]\to\mathbb{R}^{m}$
with bounded derivatives
@@
$L^{p}$ Space of random variables $\xi$ such that $\|\xi\|_{L^{p}}=\mathbb{E}[|\xi|^{p}]^{\frac{1}{p}}<\infty,$ for $p\geq 1$
$\mathbb{L}^{p}_{[0,T]}$ Space of progressively measurable processes $X$ such that
$\|X\|_{\mathbb{L}^{p}_{[0,T]}}=\bigl(\int_{0}^{T}\mathbb{E}[|X_{t}|^{p}]dt\bigr)^{\frac{1}{p}}<\infty$, for $p\geq 1$
$\mathcal{S}^{p}_{[0,T]}$ Space of progressively measurable processes $X$ such that
$\|X\|_{\mathcal{S}^{p}_{[0,T]}}=\mathbb{E}[\sup_{t\in[0,T]}|X_{t}|^{p}]^{\frac{1}{p}}<\infty,$ for $p\geq 1$
$\mathcal{S}^{\infty}_{[0,T]}$ Space of progressively measurable processes $X$ such that
$\|X\|_{\mathcal{S}^{\infty}_{[0,T]}}=\operatorname*{ess\,sup}_{\omega\in\Omega}\sup_{0\leq t\leq T}|X_{t}|<\infty$
@@
$\mathscr{L}$ Lifting map (19), $\mathbb{R}^{(n+1)\times K}\ni a\mapsto\mathscr{L}a:=a_{0}g_{0}+\cdots+a_{n}g_{n}\in L^{\infty}([0,T])$,
with $g_{0},\cdots,g_{n}$ being linearly independent vectors of $C([0,T],\mathbb{R})$
@@ 2 McKean-Vlasov SDEs with separable coefficients
Let $T>0$ be a fixed time horizon. All the stochastic differential equations in this article are considered up to time $T$. We consider the class of $\mathbb{R}^{d}$-valued MV-SDEs in (1), where the interaction through the state is separated from the interaction through the law. Their dynamics can be compactly written as
@@
for measurable maps $\alpha:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{K\times d}$, $\beta:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{K\times d\times q}$, $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{K}$, $\xi$ a random variable, and with $W$ being a $q$-dimensional Brownian motion (column vector). The full probabilistic framework can be found in the Appendix.
@@
The above result follows from [9, Theorem 4.21 (p.236)]). In particular, the fact that $\bar{\gamma}$ is bounded and continuous is straightforward as $\varphi$ is bounded and continuous.
@@
Concerning the well-posedness of MV-SDEs and the moment estimates of its solutions, Lipschitz continuity of the coefficients is a rather standard assumption. The literature on MV-SDEs also contains results for less regular coefficients (see [16] and the references therein), but these typically require additional structural assumptions, e.g. ellipticity conditions for the diffusion coefficient. Also, we rely on the Lipschitz continuity of $\alpha(t,\cdot)$ and $\beta(t,\cdot)$ in the convergence analysis of the SGD algorithm (see Section 1 below). Thus, we prefer to introduce now this regularity requirement.
@@
$0,T]\times\mathbb{R}^{d}\ni(t,x)\mapsto\big(\mathscr{K}^{m}u\big)(t,x)=\underbrace{\mathscr{K}\big(\cdots\mathscr{K}(\mathscr{K}u)\cdots\big)}_{m\text{ times}}(t,x).$ (6)
@@ 2.1 Fixed-point equation for $\bar{\gamma}$
$F:C\big([0,T],\mathbb{R}^{K}\big)\to\mathbb{R},\qquad F(\gamma):=\|\Psi(\gamma)-\gamma\|_{L^{2}_{([0,T])}}.$ (12)
@@ 2.2 Projection on a finite dimensional domain
Finally note that, in particular, the distance $d_{n}(\gamma)$ and the projection $p_{n}(\gamma)$ (see Remark 2.15) are the same for all these choices of basis, for they all generate the same finite dimensional space $S_{n}=\mathcal{P}_{n}(\mathcal{R}^{K})$. In this case, $p_{n}$ in (20) is called the polynomial of best approximation of $\gamma$ (see [34, Chapter 5]).
@@
Since the maps $\gamma$ are bounded by $\|\varphi\|_{\infty}$, we introduce the auxiliary map ${\bf h}:\mathbb{R}^{K}\to\mathbb{R}^{K}$ so that we can ensure that ${\bf h}(\mathscr{L}a)$ is also bounded. Concretely, let $\bf h$} be a smooth bounded function $\mathbb{R}^{K}\ni x\mapsto{\bf h}(x)=\big({\bf h}_{1}(x_{1}),\cdots,{\bf h}_{K}(x_{K})\big)$ defined as
@@
The role of map $\mathbf{h}$ is to ensure that the estimate of $\bar{\gamma}$, defined in (13), remains bounded when computing the loss function $F$. This is reflected when we compute $F^{2}({\bf h}\circ\mathscr{L}a)$ in (28). Moreover, the function $H$ penalizes the coefficient vector $a$ when the norm of our approximation $\mathcal{L}a$ grows beyond $\|\varphi\|_{\infty}+d_{n}(\bar{\gamma}).$ This is reflected in the definition of $H$ in (25) and also motivated in Remark 2.19.
@@
The boundedness of $\varphi$ implies that $F^{2}({\bf h}\circ\mathscr{L}a)$ is bounded. Thus, by (25), $G(a)$ tends to $\infty$ as $|a|\to\infty$. On the other hand, Lemma 2.13 together with the continuity of $\bf h$} and $H$ imply that $G$ is continuous on $\mathbb{R}^{(n+1)K}$. Therefore, the fact that $G$ has at least a minimum $\bar{a}_{n}$ on $\mathbb{R}^{(n+1)K}$ stems from Weierstrass’ extreme value theorem.
@@
A theorem of D. Jackson (see [49, Section 5.1.5 (p.261)]) then yields
@@ 3.1 Some preliminaries
    Let now $W$, $\tilde{W}$ be two independent Brownian motions and $\xi$, $\tilde{\xi}$ be two identically distributed independent random variables, which are also independent of (W,W~)($\tilde{W}$). Assuming weak-uniqueness for the solutions to (38) we have
@@ 3.2 Algorithm
    • $(W_{m})_{m\in\mathbb{N}}$ and $(\tilde{W}_{m})_{m\in\mathbb{N}}$ are sequences of independent $q$-dimensional standard Brownian motions adapted to $(\bar{\mathcal{F}}_{m})_{m\in\mathbb{N}}$. Namely, for any $m\in\mathbb{N}$, $W_{m}=(W_{m,t})_{t\in[0,T]}$ is a $q$-dimensional standard Brownian motion such that $\mathcal{F}^{W_{m}}_{T}\subset\bar{\mathcal{F}}_{m}$, and the same holds for $\tilde{W}_{m}$;
@@
Input : Maximum number of iterations $m_{\max}$, learning rates $(\eta_{m})_{m\in\mathbb{N}}$;